40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 9 12:35:47 2017, gautam, Update, ALS, Fiber ALS IMG_3942.JPG
    Reply  Mon Aug 14 16:24:09 2017, gautam, Update, ALS, Fiber ALS IMG_7471.JPGIMG_7472.JPGIMG_7473.JPGIMG_7474.JPG
       Reply  Wed Aug 16 20:24:23 2017, gautam, Update, ALS, Fiber ALS IMG_7476.JPGIMG_7477.JPGIMG_7478.JPGIMG_7479.JPG
          Reply  Wed Aug 23 17:22:36 2017, gautam, Update, ALS, Fiber ALS - reinstalled 
             Reply  Fri Aug 25 15:54:14 2017, gautam, Update, ALS, Fiber ALS noise measurement IMG_7497.JPGFOL_schematic.pdf20170825_IR_ALS.pdf
                Reply  Fri Aug 25 17:11:07 2017, rana, Update, ALS, Fiber ALS noise measurement 
                   Reply  Tue Sep 26 19:10:13 2017, gautam, Update, ALS, Fiber ALS setup neatened IMG_7605.JPG
                Reply  Sun Aug 27 11:57:31 2017, rana, Update, ALS, Fiber ALS noise measurement 
                   Reply  Tue Aug 29 02:08:39 2017, gautam, Update, ALS, Fiber ALS noise measurement 
                      Reply  Fri Sep 1 19:15:40 2017, gautam, Update, ALS, Fiber ALS noise measurement IR_ALS_noise.pdf
Message ID: 13204     Entry time: Mon Aug 14 16:24:09 2017     In reply to: 13177     Reply to this: 13222
Author: gautam 
Type: Update 
Category: ALS 
Subject: Fiber ALS 

Today, I borrowed the fiber microscope from Johannes and took a look at the fibers coupled to the PDs. The PD labelled "BEAT PD AUX Y" has an end that seems scratched (Attachments #1 and #2). The scratch seems to be on (or at least very close to) the core. The other PD (Attachments #3 and #4) doesn't look very clean either, but at least the area near the core seems undamaged. The two attachments for each PD corresponds to the two available lighting settings on the fiber microscope.

I have not attempted to clean them yet, though I have also borrowed the cleaning supplies to facilitate this from Johannes. I also plan to inspect the ends of all other fiber connections before re-installing them.

Quote:

Last week, we were talking about reviving the Fiber ALS box. Right now, it's not in great shape. Some changes to be made:

  1. Supply power to the PDs (Menlo FPD310) via a power regulator board. The datasheet says the current consumption per PD is 250 mA. So we need 500mA. We have the D1000217 power regulator board available in the lab. It uses the LM2941 and LM2991 power regulator ICs, both of which are rated for 1A output current, so this seems suitable for our purposes. Thoughts?
  2. Install power decoupling capacitors on the PDs.
  3. Clean up the fiber arrangement inside the box.
  4. Install better switches, plus LED indicators.
  5. Cover the box.
  6. Install it in a better way on the PSL table. Thoughts? e.g. can we mount the unit in some electronics rack and route the fibers to the rack? Perhaps the PSL IR and one of the arm fibers are long enough, but the other arm might be tricky.

Previous elog thread about work done on this box: elog11650

 

Attachment 1: IMG_7471.JPG  1.627 MB  | Show | Hide all | Show all
Attachment 2: IMG_7472.JPG  1.150 MB  | Hide | Hide all | Show all
IMG_7472.JPG
Attachment 3: IMG_7473.JPG  1.271 MB  | Hide | Hide all | Show all
IMG_7473.JPG
Attachment 4: IMG_7474.JPG  1.132 MB  | Hide | Hide all | Show all
IMG_7474.JPG
ELOG V3.1.3-